Write an equation of the line that passes through the given point and has slope m.
(-1,6); m = 2

Answer by mananth(16946)

The slope point formula is (y-y1)=m(x-x1)
y-6=2(x-(-1)
y-6=2x+2
2x-y+8=0
y= 2x+8